OpenStudy (anonymous):

f(x)=2x(1-cos2x) , find f'(x).

OpenStudy (anonymous):

f'(x)=2(1-cos(2x))+4x*sin(2x)

OpenStudy (anonymous):

Could you show steps?

OpenStudy (anonymous):

product rule:(uv)'=u'v+uv' u=2x u'=2 v=1-cos(2x) v'=2sin(2x)

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

welcome